# Task Diversity Shortens the ICL Plateau

### v1

@@ Abstract
11footnotetext: Co-first author22footnotetext: Our Code is available at https://github.com/sehyunkwon/task-diversity-icl.
Loss of single-task      Test error/acc of single-task
Loss of multi-task      Test error/acc of multi-task
Plateau escape point
Figure 1: We train a transformer from scratch on in-context learning tasks. Single-task ICL: Training loss ( ) and test error/accuracy ( ) when each task is trained individually. The Parity task cannot be learned even after $1000$k training steps. Multi-task ICL: Training loss ( ) and test error/accuracy ( ) when all six tasks are trained simultaneously. Green lines mark the plateau escape points. Surprisingly, multi-task ICL training significantly shortens these plateaus, making training easier.
@@ 1.1 Related works
In-context learning (ICL) abilities of pretrained Large Language Models have gained significant attention since first investigated by GPT-3 (Brown et al., 2020). Following this, a large body of empirical studies have explored ICL in Large Language Models (Min et al., 2022a; b; Liu et al., 2021; Nie et al., 2022; Rubin et al., 2022; Wei et al., 2023). Given the complexity of real-world data, researchers have explored ICL in more stylized and simplified setups. Garg et al. (2022) formalized an approach to studying transformer’s performance to learn simple function class in context. Building on this work, several works have investigated the ability of models to learn stylized function classes in various settings, including boolean functions (Bhattamishra et al., 2024), regular language (Akyürek et al., 2024), as well as task mixtures (Tripuraneni et al., 2023), and the ability of Mamba (Park et al., 2024; Grazzi et al., 2024; Li et al., 2024). For more comprehensive overview, please refer to the survey by Dong et al. (2024).
@@
Multi-task training (Caruana, 1997; Baxter, 2000) is a widely used approach for model pretraining. In particular, researchers have identified that task diversity is crucial for pre-trained models to outperform in downstream tasks, alongside a large body of research across various domains: supervised learning (Tripuraneni et al., 2020; 2021; Du et al., 2021; Maurer et al., 2016; Crawshaw, 2020; Ruder, 2017), reinforcement learning (Zhang et al., 2023a; Jin et al., 2020; Hu et al., 2021; Yang et al., 2021; Collins et al., 2021; Lu et al., 2022; Cheng et al., 2022), and natural language processing (Zhang et al., 2023a; Zhao et al., 2023; Zhang et al., 2023b; Hu et al., 2020; Song et al., 2020; Zhou et al., 2019; Gunasekar et al., 2023; Sharma et al., 2023). These works mostly focused on statistical benefits, whereas our finding emphasize on the optimization benefits.
@@ 2 Experimental setup
Figure 2: To generate a training sequence, we sample the function $f\in\mathcal{F}_{1}\cup\ldots\cup\mathcal{F}_{k}$, where $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$ are different in-context function classes, sample $x_{1},\dots,x_{n}$ IID, and form $(x_{1},f(x_{1}),\cdots,x_{n-1},f(x_{n-1}),x_{n},f(x_{n}))$. We refer to the case $k=1$ as single-task ICL and $k>1$ as multi-task ICL. The sequence model is trained with autoregressive next-token prediction, i.e., the model predicts $f(x_{i})$ conditioned on $(x_{1},f(x_{1}),x_{i-1},f(x_{i-1}),\cdots,x_{i})$ for $i=1,\dots,n$.
@@
Our experimental setup follows Garg et al. (2022) and Bhattamishra et al. (2024). Consider a function class $\mathcal{F}$ with domain $\mathcal{X}$. A sequence model $M_{\theta}$ (transformer or state-space model) is trained to identify $f\in\mathcal{F}$ in context and make a prediction on the subsequent query. The training data consists of sequences of the form $P=(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{n-1},f(x_{n-1}),x_{n},f(x_{n}))$, where $f$ is sampled from a distribution $\mathcal{D}_{\mathcal{F}}$ and $x_{1},\dots,x_{n}$ are independently sampled from a distribution $\mathcal{D}_{\mathcal{X}}$. We train $M_{\theta}$ for the next token-prediction task: The loss over the sequence $P$ is given by $\frac{1}{n}\sum_{i=1}^{n}\ell(M_{\theta}(P_{i}),f(x_{i}))$, where $\ell(\cdot,\cdot)$ is an appropriate loss function and $P_{i}:=(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{i-1},f(x_{i-1}),x_{i})$ is the $i$-th prefix for $i=1,\dots,n$. In our experiments, $n=120$ is a predetermined number shared across all tasks. This procedure is illustrated on Figure 2.
@@
ICL tasks. We consider continuous ICL tasks and boolean ICL tasks. For each ICL task, an $f\in\mathcal{F}$ is chosen, where $\mathcal{F}$ is a function class. For continuous ICL tasks, $\mathcal{F}$ consists of $f\colon\mathbb{R}^{d}\to\mathbb{R}$ for $d\in\mathbb{N}$ and the probability distribution on the domain is assumed to be $\mathcal{D}_{\mathcal{X}}=\mathcal{N}(0,I_{d})$. For boolean ICL tasks, $\mathcal{F}$ consists of $f\colon\{\pm 1\}^{d}\to\{\pm 1\}$ for $d\in\mathbb{N}$ and the probability distribution on the domain is assumed to be $\mathcal{D}_{\mathcal{X}}=\mathrm{Unif\,}(\{\pm 1\}^{d})$. We consider $10$ different ICL tasks: Linear Regression, Quadratic Regression, Sparse Linear Regression, ReLU Regression, Decision Tree, Sparse Parity($2$), Sparse Parity($3$), Parity, Conjunction, and Disjunction. The function class $\mathcal{F}$ defining each of these $10$ ICL tasks is precisely stated in Section 5.
@@
Multi-task ICL training. Most prior work on ICL, such as Garg et al. (2022); von Oswald et al. (2023); Bhattamishra et al. (2024); Park et al. (2024), considers ICL training with a single function class $\mathcal{F}$. In this work, we train models to learn functions from the union of function classes $\bigcup_{m=1}^{k}\mathcal{F}_{m}$ in context, where $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$ are distinct ICL task among the $10$ that we list in Section 5. For instance, if the model is trained on the union of linear and quadratic regression tasks, then $\mathcal{F}_{1}\cup\mathcal{F}_{2}=\{f\,|\,f(x)=w^{\intercal}x\}\cup\{f\,|\,f(x)=x^{\intercal}Wx\}$. For each $m=1,\dots,k$, we sample $f\sim\mathcal{D}_{\mathcal{F}_{m}}$ and $x_{1},\dots,x_{n}\stackrel{{\scriptstyle\text{IID}}}{{\sim}}\mathcal{D}_{\mathcal{X}_{m}}$ and form the sequence $(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{n},f(x_{n}))$. This sampling process is repeated $B$ times for each $m=1,\dots,k$, making $kB$ the total batch size. To balance the loss scales across different tasks, we normalize each task’s loss with constant factors $c_{1},\dots,c_{k}$ (further discussed in Section 4.1) so that the training loss stabilizes around $1$ during the plateaus. Appendix A provides further experimental details. In expectation, we minimize the loss
@@ 3 Task diversity shortens ICL plateaus
We provide comprehensive experimental evidence to support this claim. Table 1 summarizes our experimental results on transformers. Figure 3 shows the results of state space models, specifically Mamba (Gu & Dao, 2024) and Hyena (Poli et al., 2023). Across the hundreds of task combinations in these setups, we consistently observed that task diversity enables training to escape plateaus more quickly. Figure 7 shows results in natural language ICL tasks reinforcing our claim. Refer to Appendix A for experimental details and Appendix B for additional tables.
However, not all ICL tasks mutually reduce the duration of plateaus. For instance, we consider Regbench task (Akyürek et al., 2024), which is generated by a random automata. Our experiments in Appendix C.2 show that combining Regbench with the ICL tasks we consider does not reduce the plateau of Regbench, but does reduce the plateau of the other ICL tasks.
@@
Figure 3: Mamba (First row) and Hyena (Second row). Red lines and Blue lines respectively represent the loss dynamics of single-task training and multi-task training. (Left): Quadratic Regression$+$Linear Regression compared to Quadratic Regression. (Middle): Sparse Parity(2)$+$Linear Regression compared to Sparse Parity(2). (Right): Sparse Parity(3)$+$Conjunction compared to Sparse Parity(3).
Boolean Tasks Continuous Tasks
Number of tasks Sparse Parity(2) Sparse Parity(3) Conjunction Disjunction Linear Regression Quadratic Regression ReLU Regression Sparse Linear Regression Decision Tree
1 $>1000$k $>1000$k $2.4$k ($2.7$k) $2.5$k ($2.7$k) $9.0$k ($12.1$k) $13.1$k ($14.2$k) $3.0$k ($3.7$k) $16.5$k ($18.4$k) $4.3$k ($16.0$k)
2 $12.2$k ($17.2$k) $22.4$k ($26.7$k) $1.6$k ($1.9$k) $1.3$k ($1.6$k) $2.7$k ($3.4$k) $5.5$k ($7.5$k) $2.2$k ($2.8$k) $3.1$k ($3.8$k) $5.9$k ($19.3$k)
4 $2.1$k ($2.4$k) $2.9$k ($3.9$k) $1.5$k ($3.0$k) $1.2$k ($1.5$k) $1.4$k ($2.1$k) $2.1$k ($4.5$k) $1.8$k ($2.5$k) $1.4$k ($2.1$k) $2.5$k ($18.0$k)
8 $1.5$k ($2.0$k) $2.1$k ($2.4$k) $0.9$k ($1.5$k) $1.0$k ($1.5$k) $0.9$k ($2.0$k) $1.7$k ($5.4$k) $0.9$k ($2.0$k) $0.9$k ($1.9$k) $2.0$k ($20.0$k)
9 $1.6$k ($1.8$k) $2.2$k ($2.3$k) $0.9$k ($1.5$k) $0.9$k ($1.4$k) $0.8$k ($1.8$k) $1.6$k ($6.6$k) $0.9$k ($1.9$k) $0.8$k ($1.8$k) $1.8$k ($22.0$k)
Table 1: Task diversity shortens the ICL plateau. We train a transformer with various combinations of $9$ different tasks ($d=10$). For each run, we report two metrics: the time to escape the plateau and the time to complete the learning of the task (written in parentheses). The rows correspond to the number of tasks trained together, and each entry in the table corresponds to the average time across the training runs that include the given task. For example, the entry at (Number of tasks = 4, Conjunction) shows the average of $\binom{8}{3}$ results. We find that multi-task training shortens the ICL plateau. Precise details are provided in Appendix A.
Figure 4: Complexity model of multi-task training
@@ 4.1 Plateau is task-wise no-context learning
In other words, the model identifies the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ and then applies the optimal no-context function corresponding to $\mathcal{F}_{m}$. The in-context demonstrations $(x_{1},f(x_{1}),\dots,x_{n-1},f(x_{n-1}))$ are used to determine the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ but not to determine the specific $f\in\mathcal{F}_{m}$. This claim can be verified by measuring the error between $M_{\theta}(P_{n})$ and $g^{\star}_{\mathcal{F}}$, which we plot in Figure 5. Revisiting the Sparse Parity(2) task, the accuracy during plateau is indeed attributed to $0.55=\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}[\mathbf{1}(g_{\mathcal{F}}^{\star}(x)=f(x))]$. Appendix F provides further details on no-context learning regime.
@@ 4.2.1 Checkpoint experiment
The findings, summarized in Figure 6, indicate that the checkpoint model transferred from task A quickly learns task B with a shortened plateau. This implies that (1) task A and task B share a common structure and that (2) the plateau arises from the difficulty of learning the common structure. We conduct an analogous experiment on natural language ICL tasks and observe similar results, as shown in Figure 8 of the appendix. Further details are presented in Appendix D.

### v2

# Task Diversity Shortens the In-Context Learning Plateau

@@ Abstract
Machine Learning, ICML
Figure 1: We train a transformer from scratch on in-context learning tasks. Single-task ICL: Training loss ( ) and test error/accuracy ( ) when each task is trained individually. The Parity task cannot be learned even after $1000$k training steps. Multi-task ICL: Training loss ( ) and test error/accuracy ( ) when all five tasks are trained simultaneously. Green lines mark the plateau escape points. Surprisingly, multi-task ICL training significantly shortens these plateaus, making training easier.
@@ 1.1 Related works
In-context learning (ICL) abilities of pretrained Large Language Models have gained significant attention since first investigated by GPT-3 (Brown et al., 2020). Following this, a large body of empirical studies have explored ICL in Large Language Models (Min et al., 2022a, b; Liu et al., 2021; Nie et al., 2022; Rubin et al., 2022; Wei et al., 2023). Given the complexity of real-world data, researchers have explored ICL in more stylized and simplified setups. Garg et al. (2022) formalized an approach to studying transformer’s performance to learn simple function class in context. Building on this work, several works have investigated the ability of models to learn stylized function classes in various settings, including boolean functions (Bhattamishra et al., 2024), regular language (Akyürek et al., 2024), as well as task mixtures (Tripuraneni et al., 2023), and the ability of Mamba (Park et al., 2024; Grazzi et al., 2024; Li et al., 2024). For more comprehensive overview, please refer to the survey by Dong et al. (2024).
@@
Multi-task training (Caruana, 1997; Baxter, 2000) is a widely used approach for model pretraining. In particular, researchers have identified that task diversity is crucial for pre-trained models to outperform in downstream tasks, alongside a large body of research across various domains: supervised learning (Tripuraneni et al., 2020, 2021; Du et al., 2021; Maurer et al., 2016; Crawshaw, 2020; Ruder, 2017), reinforcement learning (Zhang et al., 2023a; Jin et al., 2020; Hu et al., 2021; Yang et al., 2021; Collins et al., 2021; Lu et al., 2022; Cheng et al., 2022), and natural language processing (Zhang et al., 2023a; Zhao et al., 2023; Zhang et al., 2023b; Hu et al., 2020; Song et al., 2020; Zhou et al., 2019; Gunasekar et al., 2023; Sharma et al., 2023). These works mostly focused on statistical benefits, whereas our finding emphasize on the optimization benefits.
Figure 2: To generate a training sequence, we sample the function $f\in\mathcal{F}_{1}\cup\dots\cup\mathcal{F}_{k}$, where $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$ are different in-context function classes, sample $x_{1},\dots,x_{n}$ IID, and form $(x_{1},f(x_{1}),\dots,x_{n-1},f(x_{n-1}),x_{n},f(x_{n}))$. We refer to the case $k=1$ as single-task ICL and $k>1$ as multi-task ICL. The sequence model is trained with autoregressive next-token prediction, i.e., the model predicts $f(x_{i})$ conditioned on $(x_{1},f(x_{1}),\dots,x_{i-1},f(x_{i-1}),x_{i})$ for $i=1,\dots,n$.
@@ 2 Experimental setup
Our experimental setup follows Garg et al. (2022) and Bhattamishra et al. (2024). Consider a function class $\mathcal{F}$ with domain $\mathcal{X}$. A sequence model $M_{\theta}$ (transformer or state-space model) is trained to identify $f\in\mathcal{F}$ in context and make a prediction on the subsequent query. The training data consists of sequences of the form $P=(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{n-1},f(x_{n-1}),x_{n},f(x_{n}))$, where $f$ is sampled from a distribution $\mathcal{D}_{\mathcal{F}}$ and $x_{1},\dots,x_{n}$ are independently sampled from a distribution $\mathcal{D}_{\mathcal{X}}$. We train $M_{\theta}$ for the next token-prediction task: The loss over the sequence $P$ is given by $\frac{1}{n}\sum_{i=1}^{n}\ell(M_{\theta}(P_{i}),f(x_{i}))$, where $\ell(\cdot,\cdot)$ is an appropriate loss function and $P_{i}=(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{i-1},f(x_{i-1}),x_{i})$ is the $i$-th prefix for $i=1,\dots,n$. In our experiments, $n=120$ is a predetermined number shared across all tasks. This procedure is illustrated in Figure 2.
@@
ICL tasks. We consider continuous ICL tasks and boolean ICL tasks. For each ICL task, an $f\in\mathcal{F}$ is chosen, where $\mathcal{F}$ is a function class. For continuous ICL tasks, $\mathcal{F}$ consists of $f\colon\mathbb{R}^{d}\to\mathbb{R}$ for $d\in\mathbb{N}$ and the probability distribution on the domain is assumed to be $\mathcal{D}_{\mathcal{X}}=\mathcal{N}(0,I_{d})$. For boolean ICL tasks, $\mathcal{F}$ consists of $f\colon\{\pm 1\}^{d}\to\{\pm 1\}$ for $d\in\mathbb{N}$ and the probability distribution on the domain is assumed to be $\mathcal{D}_{\mathcal{X}}=\mathrm{Unif\,}(\{\pm 1\}^{d})$. We consider $7$ different ICL tasks: Linear Regression, Quadratic Regression, Sparse Linear Regression, LeakyReLU Regression, Sparse Parity($2$), Sparse Parity($3$), Parity. The function class $\mathcal{F}$ defining each of these $7$ ICL tasks is precisely stated in Section 5.
@@
Multi-task ICL training. Most prior work on ICL, such as Garg et al. (2022); von Oswald et al. (2023); Bhattamishra et al. (2024); Park et al. (2024), considers ICL training with a single function class $\mathcal{F}$. In this work, we train models to learn functions from the union of function classes $\bigcup_{m=1}^{k}\mathcal{F}_{m}$ in context, where $\mathcal{F}_{1},\dots,\mathcal{F}_{k}$ are distinct ICL task among the $10$ that we list in Section 5. For instance, if the model is trained on the union of linear and quadratic regression tasks, then $\mathcal{F}_{1}\cup\mathcal{F}_{2}=\{f\,|\,f(x)=w^{\intercal}x\}\cup\{f\,|\,f(x)=x^{\intercal}Wx\}$. For each $m=1,\dots,k$, we sample $f\sim\mathcal{D}_{\mathcal{F}_{m}}$ and $x_{1},\dots,x_{n}\stackrel{{\scriptstyle\text{IID}}}{{\sim}}\mathcal{D}_{\mathcal{X}_{m}}$ and form the sequence $(x_{1},f(x_{1}),x_{2},f(x_{2}),\dots,x_{n},f(x_{n}))$. This sampling process is repeated $B/k$ times for each $m=1,\dots,k$, making $B$ the total batch size. This choice of batch size ensures a fair comparison between single- and multi-task training, as the number of tokens during training is exactly proportional to the iteration count. Additionally, we consider a multi-task mixture with uneven probabilities to further support the generality of our claim. To balance the loss scales across different tasks, we normalize each task’s loss with constant factors $c_{1},\dots,c_{k}$ (further discussed in Section 4.1) so that the training loss stabilizes around $1$ during the plateaus. Appendix A provides further experimental details. In expectation, we minimize the loss
@@ 3 Task diversity shortens ICL plateaus
Table 1: Task diversity shortens the ICL plateau. We train a transformer with various combinations of $6$ different tasks ($d=10$). For each run, we report two metrics: the time to escape the plateau and the time to complete the learning of the task (written in parentheses). The rows correspond to the number of tasks trained together, and each entry in the table corresponds to the average time across the training runs that include the given task. For example, the entry at (Number of tasks = 3, Sparse Parity(2)) shows the average of $\binom{6}{3}$ results. We find that multi-task training shortens the ICL plateau. Precise details are provided in Appendix A.
Boolean Tasks Continuous Tasks
Number of tasks Sparse Parity(2) Sparse Parity(3) Linear Regression Quadratic Regression LeakyReLU Regression Sparse Linear Regression
1 $>1000$k $>1000$k $9.0$k ($12.1$k) $13.1$k ($14.2$k) $7.4$k ($8.9$k) $16.5$k ($18.4$k)
2 $16.1$k ($26.5$k) $19.9$k ($32.1$k) $6.1$k ($7.6$k) $7.6$k ($9.5$k) $3.7$k ($5.0$k) $4.7$k ($6.6$k)
3 $5.3$k ($5.8$k) $7.1$k ($11.3$k) $3.0$k ($3.9$k) $3.9$k ($5.8$k) $2.3$k ($3.0$k) $3.7$k ($4.7$k)
4 $4.8$k ($5.7$k) $6.2$k ($10.1$k) $3.7$k ($4.7$k) $3.1$k ($5.5$k) $2.6$k ($3.4$k) $4.3$k ($5.3$k)
5 $3.2$k ($3.9$k) $4.6$k ($7.9$k) $3.0$k ($3.9$k) $3.4$k ($5.9$k) $2.7$k ($3.6$k) $3.0$k ($3.9$k)
6 $2.3$k ($3.0$k) $3.2$k ($5.0$k) $1.9$k ($2.4$k) $2.4$k ($4.8$k) $1.9$k ($2.4$k) $1.9$k ($2.4$k)
@@
Generality of our claim.
We provide comprehensive experimental evidence to support this claim. Table 1 summarizes our experimental results on transformers. Figure 3 shows the results of state space models, specifically Mamba (Gu & Dao, 2024) and Hyena (Poli et al., 2023). Across the hundreds of task combinations in these setups, we consistently observed that task diversity allows training to escape plateaus more quickly. Table 2 presents results for multi-task learning with uneven probabilities of mixture. Figure 7 shows results in natural language ICL tasks supporting our claim. Refer to Appendix A for experimental details.
However, not all ICL tasks mutually reduce the duration of plateaus. For instance, we consider Regbench task (Akyürek et al., 2024), which is generated by a random automata. Our experiments in Appendix B.2 show that combining Regbench with the ICL tasks we consider does not reduce the plateau of Regbench, but does reduce the plateau of the other ICL tasks.
@@
Table 2: Muti-task ICL with uneven probabilities. We conducted experiments on multi-task ICL with uneven mixture probabilities, where $5$ ICL tasks are mixed with the ratio $\left(\frac{1}{2},\frac{1}{8},\frac{1}{8},\frac{1}{8},\frac{1}{8}\right)$. Under these uneven mixtures, we consistently observed a shortened plateau.
SP(2) SP(3) LR QR Sparse LR
1 $\infty$ $\infty$ $2.8$k ($4.2$k) $10.2$k ($11.5$k) $2.3$k ($3.3$k)
5 $2.4$k ($2.9$k) $3.0$k ($6.8$k) $2.0$k ($2.5$k) $2.4$k ($4.1$k) $2.0$k ($2.5$k)
Figure 3: Mamba (First row) and Hyena (Second row). Red lines and Blue lines respectively represent the loss dynamics of single-task training and multi-task training. (Left): Quadratic Regression vs. Quadratic Regression$+$Linear Regression. (Middle): Sparse Parity(2) vs. Sparse Parity(2)$+$Linear Regression. (Right): Sparse Parity(3) vs. Sparse Parity(3)$+$Quadratic Regression.
Figure 4: Complexity model of multi-task training. The illustration demonstrates the theoretical framework and computational aspects of multi-task training.
@@ 4.1 Plateau is task-wise no-context learning
In other words, the model identifies the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ and then applies the optimal no-context function corresponding to $\mathcal{F}_{m}$. The in-context demonstrations $(x_{1},f(x_{1}),\dots,x_{n-1},f(x_{n-1}))$ are used to determine the function class $\mathcal{F}_{m}\in\{\mathcal{F}_{1},\dots,\mathcal{F}_{k}\}$ but not to determine the specific $f\in\mathcal{F}_{m}$. This claim can be verified by measuring the error between $M_{\theta}(P_{n})$ and $g^{\star}_{\mathcal{F}}$, which we plot in Figure 5. Revisiting the Sparse Parity(2) task, the accuracy during plateau is indeed attributed to $0.55=\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}[\mathbf{1}(g_{\mathcal{F}}^{\star}(x)=f(x))]$. Appendix E provides further details on no-context learning regime.
@@ 4.2.1 Checkpoint experiment
The findings, summarized in Figure 6, indicate that the checkpoint model transferred from task A quickly learns task B with a shortened plateau. This implies that (1) task A and task B share a common structure and that (2) the plateau arises from the difficulty of learning the common structure. We conduct an analogous experiment on natural language ICL tasks and observe similar results, as shown in Figure 8 of the appendix. Further details are presented in Appendix C.